OpenStudy (jessiegonzales):

MEDAL!!! ALGEBRA 2 1. What is the slope of the line that passes throughout the given points? (2,12) AND (6,11) -1/4 1/4 4 -4

rishavraj (rishavraj):

\[\frac{ y_1 - y_2 }{ x_1 - x_2 } = m\]

OpenStudy (jessiegonzales):

2-1/12-11?

rishavraj (rishavraj):

\[\frac{ 12 - 11 }{ 2 - 6 }\]

OpenStudy (jessiegonzales):

1 - -4

rishavraj (rishavraj):

u mean \[\frac{ -1 }{ 4 }\] thts correct

OpenStudy (jessiegonzales):

yes a is the answer

rishavraj (rishavraj):

yeah :))

OpenStudy (jessiegonzales):

<3